# Delving into Dynamic Scene Cue-Consistency for Robust 3D Multi-Object Tracking

### v1

@@ Abstract
3D multi-object tracking is a critical and challenging task in the field of autonomous driving. A common paradigm relies on modeling individual object motion, e.g., Kalman filters, to predict trajectories. While effective in simple scenarios, this approach often struggles in crowded environments or with inaccurate detections, as it overlooks the rich geometric relationships between objects. This highlights the need to leverage spatial cues. However, existing geometry-aware methods can be susceptible to interference from irrelevant objects, leading to ambiguous features and incorrect associations. To address this, we propose focusing on cue-consistency: identifying and matching stable spatial patterns over time. We introduce the Dynamic Scene Cue-Consistency Tracker (DSC-Track) to implement this principle. Firstly, we design a unified spatiotemporal encoder using Point Pair Features (PPF) to learn discriminative trajectory embeddings while suppressing interference. Secondly, our cue-consistency transformer module explicitly aligns consistent feature representations between historical tracks and current detections. Finally, a dynamic update mechanism preserves salient spatiotemporal information for stable online tracking. Extensive experiments on the nuScenes and Waymo Open Datasets validate the effectiveness and robustness of our approach. On the nuScenes benchmark, for instance, our method achieves state-of-the-art performance, reaching 73.2% and 70.3% AMOTA on the validation and test sets, respectively.
@@ 1 Introduction
Accurate and reliable 3D Multi-Object Tracking (MOT) is essential for ensuring the safety and reliability of autonomous driving systems. With the steady improvement of 3D object detectors (Qi et al. 2018; Zhou and Tuzel 2018; Lang et al. 2019; Yin, Zhou, and Krahenbuhl 2021), the tracking-by-detection paradigm (Zhang et al. 2021; Xinshuo Weng 2020; Zeng et al. 2021) has remained a popular choice due to its ability to leverage rich appearance, motion, and geometric information of objects.
@@
One mainstream category of methods focuses on data association using predefined motion models (Yin, Zhou, and Krahenbuhl 2021; Weng et al. 2020) (see Figure 1(Left)). Although these methods have achieved significant progress, their performance remains limited in complex scenarios due to their difficulty in effectively capturing multi-frame temporal information. To address this limitation, another line of work (Kim et al. 2022; Chu et al. 2023) enhances feature representations by fusing geometric information with spatial dependencies of objects. However, this approach introduces new challenges: on one hand, its performance is highly susceptible to inaccuracies from upstream detectors; on the other hand, the indiscriminate utilization of spatial neighborhood nodes leads to temporally inconsistent feature representations, resulting in erroneous associations.
@@
Specifically, DSC-Track consists of the following steps:  (1) A memory bank is constructed for each trajectory to store the historical information of the tracked object and the indices of spatially dependent nodes.  (2) The unified spatiotemporal encoder aggregates unique feature embeddings of each trajectory from the historical dynamic space through Point Pair Features (PPF) (Drost et al. 2010) as spatial geometric representations, significantly reducing interference from irrelevant nodes.  (3) A Transformer-based cue-consistency interaction module leverages spatial cues to extract consistent feature representations between trajectory embeddings and detections.  (4) Feature-level affinity matrices constructed from consistency features enable data association, while the spatially dependent nodes in the memory bank are updated based on the latest dependencies of detection nodes mined by the cue-consistency interaction module, achieving robust object tracking in dynamic scenes.
To validate our approach, we conducted extensive experiments on two large-scale benchmarks: the nuScenes (Caesar et al. 2020) and Waymo Open Datasets  (Sun et al. 2020). Our ablation studies confirm the effectiveness of each proposed component in improving tracking performance and reducing heuristic dependencies. The key contributions of this work are summarized as follows:
@@
  • Using CenterPoint (Yin, Zhou, and Krahenbuhl 2021) detections as input, our method establishes a new state-of-the-art on the nuScenes benchmark, achieving 73.2% and 70.3% AMOTA on the validation and test sets, respectively, and demonstrates strong, generalizable performance on the Waymo dataset, confirming the robustness of our design.
@@ 2 Related Work
The Tracking-by-Detection (TBD) paradigm is dominant in both 2D and 3D Multi-Object Tracking (MOT). Early 2D methods, such as SORT (Bewley et al. 2016), relied on motion models like the Kalman filter for state prediction. Subsequent works (Wojke, Bewley, and Paulus 2017; Zhang et al. 2021) improved upon this by incorporating appearance features and refining association strategies. This philosophy extends to 3D MOT, where methods often pair high-fidelity detectors with similar motion-based association techniques (Weng et al. 2020; Zhang et al. 2023). To capture more complex object interactions, learning-based approaches have emerged, utilizing Graph Neural Networks (GNNs) (Brasó and Leal-Taixé 2020; Zaech et al. 2022) or Transformers (Chu et al. 2023; Ding et al. 2022) to model relationships between objects. While effective, many of these methods primarily focus on individual object cues and overlook the rich structural information present in the spatial context.
@@
Leveraging spatial information is crucial for robust tracking in crowded scenes. Some methods model spatial relationships using graph structures (He et al. 2021; Kim et al. 2022) or Transformers (Chu et al. 2023; Ding et al. 2023). However, they often face a common dilemma: they either heuristically restrict the interaction range to local neighborhoods, missing important long-range dependencies, or they indiscriminately aggregate global information (Huang, Yang, and Tsai 2023), which introduces significant interference from irrelevant objects and leads to ambiguous feature representations. In contrast to these approaches, our method addresses this challenge by adaptively searching for a unique set of spatially consistent dependent nodes for each track. This allows us to aggregate highly relevant contextual and geometric cues while suppressing interference, thereby generating more discriminative features for robust association.
@@ 3.1 Preliminaries
Given a set of $N$ 3D bounding box candidates, $\mathbf{B}^{t}=\{\mathbf{b}_{j}^{t}\mid j=1,\dots,N\}$, detected at the current frame $t$, and a set of $M$ active tracks, $\mathbf{T}^{t-1}=\{\tau_{i}^{t-1}\mid i=1,\dots,M\}$, from the previous frame, our primary goal in 3D Multi-Object Tracking is to establish associations between the active tracks in $\mathbf{T}^{t-1}$ and the new detections in $\mathbf{B}^{t}$. These inputs are typically provided by a 3D object detector, such as CenterPoint (Yin, Zhou, and Krahenbuhl 2021).
@@
Each detection candidate $\mathbf{b}_{i}$ in $\mathbf{B}^{t}$ is represented by a state vector, following conventions from prior works (Zaech et al. 2022; Ding et al. 2023). Specifically, for the nuScenes dataset, this state is defined as $\mathbf{b}_{i}=[p_{i},\theta_{i},s_{i},v_{i},c_{i},o_{i}]\in\mathbb{R}^{17}$. This vector encapsulates the 3D center position $p_{i}\in\mathbb{R}^{3}$, heading angle $\theta_{i}\in\mathbb{R}$, 3D box size $s_{i}\in\mathbb{R}^{3}$, velocity $v_{i}\in\mathbb{R}^{2}$, a one-hot encoded class label $c_{i}\in\mathbb{R}^{7}$, and the detection confidence score $o_{i}\in\mathbb{R}$.
@@ 3.2 Unified Spatiotemporal Aggregation
Modeling geometric relationships from absolute object states is unreliable, as these features lack rotation-invariance and become unstable when objects maneuver (e.g., a vehicle turning). To address this, we construct a local geometric embedding using the Point Pair Feature (PPF) (Drost et al. 2010). By capturing the relative geometry—defined by the distance and inter-object angles (see Figure 3(Left))—PPF provides a robust representation that remains consistent across rotations and viewpoint changes, ensuring stable geometric modeling.
@@
This geometric matrix is then processed to create our final feature triplet. First, $\mathbf{E}_{i}$ is enhanced with sinusoidal positional encodings (Vaswani et al. 2017) and projected by an MLP to form the relative geometry embedding $\mathbf{R}_{i}\in\mathbb{R}^{k\times d}$. Concurrently, the state vectors of the reference object $\mathbf{b}_{i}$ and its neighbors $\{\mathbf{b}_{j}\}_{j\in\mathcal{K}_{i}}$ are projected by a shared MLP into contextual embeddings: the reference feature $\mathbf{f}_{i}\in\mathbb{R}^{1\times d}$ and neighbor features $\mathbf{F}_{i}\in\mathbb{R}^{k\times d}$.
@@
The core of this process is our Geometric Inject Attention (GIA) module, inspired by (Yu et al. 2023) and depicted in Figure 3 (right). The GIA module enhances standard attention by injecting geometric information into both the attention score calculation and the value aggregation. The process involves several key steps. First, we project the input triplet into components for the attention mechanism. The contextual embeddings are projected into a standard query, key, and value:
@@
where $\text{LN}(\cdot)$ is Layer Normalization (Ba, Kiros, and Hinton 2016) and FFN consists of a two-layer MLP with a residual connection.
@@
Inspired by BERT (Devlin 2018; Huang, Yang, and Tsai 2023), we introduce a learnable track token, $\mathbf{z}_{m}\in\mathbb{R}^{1\times d}$, which acts as a summary for the entire trajectory. This token is prepended to the historical sequence, and the combined input is fed into a self-attention module. The query, key, and value matrices are computed as follows:
@@ 3.3 Cue-Consistent Attention Module
First, we independently enhance the features of tracks and detections. For the $M$ input track features, aggregated in a matrix $\hat{\mathbf{Z}}_{m}\in\mathbb{R}^{M\times d}$, we use a vanilla self-attention layer (Vaswani et al. 2017) to mine intra-feature relationships, producing globally-aware features $\tilde{\mathbf{Z}}_{m}\in\mathbb{R}^{M\times d}$.
@@
Cross-attention is a typical module for exchanging information between two feature sets (Huang et al. 2021; Wang and Solomon 2019; Sun et al. 2021). Inspired by (Chen et al. 2024), we design a cue-consistent cross-attention module to model the consistency between $\tilde{\mathbf{Z}}_{m}$ and $\tilde{\mathbf{Z}}_{B}$. For clarity, we demonstrate the process for updating detection features by attending to tracks; the reverse process is symmetric.
@@
It is important to note that computing cue-consistent cross-attention for all pairs is computationally demanding. To reduce this burden, we utilize prior knowledge from online MOT. We first predict the potential locations of historical trajectories in the current frame by applying an FFN: $p_{m}=\text{FFN}(\{\hat{\mathbf{Z}}_{m}\})$. Interactions are then established only between objects of the same class, and are further pruned based on a class-specific distance threshold. This threshold is calculated from dataset statistics of maximum speed per class, following (Zaech et al. 2022).
@@ 3.4 Feature Matching and Update
The model is trained with a composite loss. For association, we use the Focal Loss (Lin et al. 2017) ($\mathcal{L}_{a}$) with $\alpha=-1,\gamma=1$. For position regression, we use a smooth $\ell_{1}$ loss ($\mathcal{L}_{p}$). The total loss is:
@@ 4.1 Experiment Setup
We conduct experiments on two large-scale autonomous driving benchmarks. The Waymo Open Dataset (Sun et al. 2020) provides 1150 sequences (798 training, 202 validation, 150 testing) of 20-second driving data with 3D labels for three classes: Vehicle, Pedestrian, and Cyclist. The nuScenes dataset (Caesar et al. 2020), our primary benchmark, comprises 1000 sequences of similar length. It is distinguished by a richer sensor suite (32-beam LiDAR, RADAR, six cameras) and provides 3D tracking annotations at 2Hz for seven object classes, making it ideal for leveraging multi-modal data.
@@
We follow the official nuScenes tracking benchmark protocol for evaluation. The primary metrics are Average Multi-Object Tracking Accuracy (AMOTA) and Average Multi-Object Tracking Precision (AMOTP) (Weng et al. 2020). We also report secondary metrics from CLEAR MOT (Bernardin and Stiefelhagen 2008), including Multi-Object Tracking Accuracy (MOTA), ID Switches (IDS).
@@
We construct the training dataset using outputs from CenterPoint (Yin, Zhou, and Krahenbuhl 2021).Following 3DMOTFormer (Ding et al. 2023), We trained our tracker for 16 epochs on randomly sample mini-sequences of length TT=10. For the memory buffer, we set the maximum temporal length $T_{max}$=7 and select the top $k=3$ spatially dependent objects to capture the historical information for each tracklet. In the cue-consistent attention module, we set top $k=3$ to perform the cue-consistent cross attention. In the optimizing process, we use AdamW (Loshchilov 2017) optimizer with an initial learning rate of $2e^{-4}$ and cosine decay learning rate strategy with power set to 0.8. All experiments are trained on eight NVIDIA 4090 GPUs. Each GPU holds one batch element.
@@ 4.2 Main Results
To validate our approach, we compare DSC-Track against state-of-the-art methods on the nuScenes and Waymo benchmarks. On nuScenes, we compare against two categories of baselines: methods leveraging additional 3D appearance cues like ShaSTA (Sadjadpour et al. 2023) and NEBP (Liang and Meyer 2023), and those without, including strong competitors like 3DMOTFormer (Ding et al. 2023), PolarMOT (Kim et al. 2022), ImmortalTracker (Wang et al. 2021), SimpleTrack (Pang, Li, and Wang 2022), and CBMOT (Benbarka, Schröder, and Zell 2021). On Waymo, we compare against recent strong methods such as SpOT (Stearns et al. 2022) and TrajectoryFormer (Chen et al. 2023). For fair comparison, all methods are built upon detections from the widely-used CenterPoint (Yin, Zhou, and Krahenbuhl 2021) detector.
@@
This strong performance is mirrored on the nuScenes validation set (Table 2), where DSC-Track again sets a new state-of-the-art with 73.2% AMOTA, surpassing 3DMOTFormer by 2.0%. To evaluate generalizability, we also test on the Waymo Open Dataset (Table 3), where our method excels at tracking stability, securing the best or tied-for-best IDS scores across all major categories: Vehicle, Pedestrian, and Cyclist. This superior identity preservation on a different large-scale dataset validates the effectiveness and generalizability of our cue-consistency mechanism.

### v2

@@ Abstract
3D multi-object tracking is a critical and challenging task in the field of autonomous driving. A common paradigm relies on modeling individual object motion, e.g., Kalman filters, to predict trajectories. While effective in simple scenarios, this approach often struggles in crowded environments or with inaccurate detections, as it overlooks the rich geometric relationships between objects. This highlights the need to leverage spatial cues. However, existing geometry-aware methods can be susceptible to interference from irrelevant objects, leading to ambiguous features and incorrect associations. To address this, we propose focusing on cue-consistency: identifying and matching stable spatial patterns over time. We introduce the Dynamic Scene Cue-Consistency Tracker (DSC-Track) to implement this principle. Firstly, we design a unified spatiotemporal encoder using Point Pair Features (PPF) to learn discriminative trajectory embeddings while suppressing interference. Secondly, our cue-consistency transformer module explicitly aligns consistent feature representations between historical tracks and current detections. Finally, a dynamic update mechanism preserves salient spatiotemporal information for stable online tracking. Extensive experiments on the nuScenes and Waymo Open Datasets validate the effectiveness and robustness of our approach. On the nuScenes benchmark, for instance, our method achieves state-of-the-art performance, reaching 73.2% and 70.3% AMOTA on the validation and test sets, respectively. Code Repository: https://github.com/zhn12343333/DSCTrack
@@ 1 Introduction
Accurate and reliable 3D Multi-Object Tracking (MOT) is essential for ensuring the safety and reliability of autonomous driving systems. With the steady improvement of 3D object detectors (qi2018frustum; zhou2018voxelnet; lang2019pointpillars; yin2021center), the tracking-by-detection paradigm (Zhang2021ByteTrackMT; Weng2020GNN3DMOT; Zeng2021Alphatrack) has remained a popular choice due to its ability to leverage rich appearance, motion, and geometric information of objects.
@@
One mainstream category of methods focuses on data association using predefined motion models (yin2021center; Weng20193DMT) (see Figure 1(Left)). Although these methods have achieved significant progress, their performance remains limited in complex scenarios due to their difficulty in effectively capturing multi-frame temporal information. To address this limitation, another line of work (kim2022polarmot; chu2023transmot) enhances feature representations by fusing geometric information with spatial dependencies of objects. However, this approach introduces new challenges: on one hand, its performance is highly susceptible to inaccuracies from upstream detectors; on the other hand, the indiscriminate utilization of spatial neighborhood nodes leads to temporally inconsistent feature representations, resulting in erroneous associations.
@@
Specifically, DSC-Track consists of the following steps:  (1) A memory bank is constructed for each trajectory to store the historical information of the tracked object and the indices of spatially dependent nodes.  (2) The unified spatiotemporal encoder aggregates unique feature embeddings of each trajectory from the historical dynamic space through Point Pair Features (PPF) (drost2010model) as spatial geometric representations, significantly reducing interference from irrelevant nodes.  (3) A Transformer-based cue-consistency interaction module leverages spatial cues to extract consistent feature representations between trajectory embeddings and detections.  (4) Feature-level affinity matrices constructed from consistency features enable data association, while the spatially dependent nodes in the memory bank are updated based on the latest dependencies of detection nodes mined by the cue-consistency interaction module, achieving robust object tracking in dynamic scenes.
To validate our approach, we conducted extensive experiments on two large-scale benchmarks: the nuScenes (Caesar2019nuScenesAM) and Waymo Open Datasets  (sun2020scalability). Our ablation studies confirm the effectiveness of each proposed component in improving tracking performance and reducing heuristic dependencies. The key contributions of this work are summarized as follows:
@@
  • Using CenterPoint (yin2021center) detections as input, our method establishes a new state-of-the-art on the nuScenes benchmark, achieving 73.2% and 70.3% AMOTA on the validation and test sets, respectively, and demonstrates strong, generalizable performance on the Waymo dataset, confirming the robustness of our design.
@@ 2 Related Work
The Tracking-by-Detection (TBD) paradigm dominates Multi-Object Tracking (MOT). Early 2D methods evolved from simple motion models like SORT (Bewley2016SimpleOA) to more advanced techniques incorporating appearance features and refined association strategies (Wojke2017SimpleOA; Zhang2021ByteTrackMT). This philosophy extends to 3D MOT, where methods often pair high-fidelity detectors with similar association techniques (Weng20193DMT; zhang2023bytetrackv2). More recent learning-based approaches utilize Graph Neural Networks (GNNs) (braso2020learning; zaech2022learnable) or Transformers (chu2023transmot; Ding2022End2End) to model relationships, yet many still primarily focus on individual object cues, overlooking rich structural information in the spatial context.
@@
Leveraging spatial information is crucial for robust tracking. While some methods model spatial relationships using graphs (he2021learnable; kim2022polarmot) or Transformers (chu2023transmot; Ding2023MOTFormer), they often face a dilemma: either restricting interactions to local neighborhoods or indiscriminately aggregating global information, which introduces interference from irrelevant objects. In contrast, our method addresses this by adaptively searching for a unique set of spatially consistent dependent nodes for each track. This allows aggregation of highly relevant contextual cues while suppressing interference for more robust association.
@@ 3.1 Preliminaries
Given a set of $N$ 3D bounding box candidates, $\mathbf{B}^{t}=\{\mathbf{b}_{j}^{t}\mid j=1,\dots,N\}$, detected at the current frame $t$, and a set of $M$ active tracks, $\mathbf{T}^{t-1}=\{\tau_{i}^{t-1}\mid i=1,\dots,M\}$, from the previous frame, our primary goal in 3D Multi-Object Tracking is to establish associations between the active tracks in $\mathbf{T}^{t-1}$ and the new detections in $\mathbf{B}^{t}$. These inputs are typically provided by a 3D object detector, such as CenterPoint (yin2021center).
@@
Each detection candidate $\mathbf{b}_{i}$ in $\mathbf{B}^{t}$ is represented by a state vector, following conventions from prior works (zaech2022learnable; Ding2023MOTFormer). Specifically, for the nuScenes dataset, this state is defined as $\mathbf{b}_{i}=[p_{i},\theta_{i},s_{i},v_{i},c_{i},o_{i}]\in\mathbb{R}^{17}$. This vector encapsulates the 3D center position $p_{i}\in\mathbb{R}^{3}$, heading angle $\theta_{i}\in\mathbb{R}$, 3D box size $s_{i}\in\mathbb{R}^{3}$, velocity $v_{i}\in\mathbb{R}^{2}$, a one-hot encoded class label $c_{i}\in\mathbb{R}^{7}$, and the detection confidence score $o_{i}\in\mathbb{R}$.
@@ 3.2 Unified Spatiotemporal Aggregation
Modeling geometric relationships from absolute object states is unreliable, as these features lack rotation-invariance and become unstable when objects maneuver (e.g., a vehicle turning). To address this, we construct a local geometric embedding using the Point Pair Feature (PPF) (drost2010model). By capturing the relative geometry—defined by the distance and inter-object angles (see Figure 3(Left))—PPF provides a robust representation that remains consistent across rotations and viewpoint changes, ensuring stable geometric modeling.
@@
This geometric matrix is then processed to create our final feature triplet. First, $\mathbf{E}_{i}$ is enhanced with sinusoidal positional encodings (Vaswani2017attention) and projected by an MLP to form the relative geometry embedding $\mathbf{R}_{i}\in\mathbb{R}^{k\times d}$. Concurrently, the state vectors of the reference object $\mathbf{b}_{i}$ and its neighbors $\{\mathbf{b}_{j}\}_{j\in\mathcal{K}_{i}}$ are projected by a shared MLP into contextual embeddings: the reference feature $\mathbf{f}_{i}\in\mathbb{R}^{1\times d}$ and neighbor features $\mathbf{F}_{i}\in\mathbb{R}^{k\times d}$.
@@
The core of this process is our Geometric Inject Attention (GIA) module, inspired by (yu2023rotation) and depicted in Figure 3 (right). The GIA module enhances standard attention by injecting geometric information into both the attention score calculation and the value aggregation. The process involves several key steps. First, we project the input triplet into components for the attention mechanism. The contextual embeddings are projected into a standard query, key, and value:
@@
where $\text{LN}(\cdot)$ is Layer Normalization (Ba2016LayerN) and FFN consists of a two-layer MLP with a residual connection.
@@
Inspired by BERT (devlin2018bert; huang2023delving), we introduce a learnable track token, $\mathbf{z}_{m}\in\mathbb{R}^{1\times d}$, which acts as a summary for the entire trajectory. This token is prepended to the historical sequence, and the combined input is fed into a self-attention module. The query, key, and value matrices are computed as follows:
@@ 3.3 Cue-Consistent Attention Module
First, we independently enhance the features of tracks and detections. For the $M$ input track features, aggregated in a matrix $\hat{\mathbf{Z}}_{m}\in\mathbb{R}^{M\times d}$, we use a vanilla self-attention layer (Vaswani2017attention) to mine intra-feature relationships, producing globally-aware features $\tilde{\mathbf{Z}}_{m}\in\mathbb{R}^{M\times d}$.
@@
Cross-attention is a typical module for exchanging information between two feature sets (huang2021predator; wang2019deep; sun2021loftr). Inspired by (chen2024dynamic), we design a cue-consistent cross-attention module to model the consistency between $\tilde{\mathbf{Z}}_{m}$ and $\tilde{\mathbf{Z}}_{B}$. For clarity, we demonstrate the process for updating detection features by attending to tracks; the reverse process is symmetric.
@@
It is important to note that computing cue-consistent cross-attention for all pairs is computationally demanding. To reduce this burden, we utilize prior knowledge from online MOT. We first predict the potential locations of historical trajectories in the current frame by applying an FFN: $p_{m}=\text{FFN}(\{\hat{\mathbf{Z}}_{m}\})$. Interactions are then established only between objects of the same class, and are further pruned based on a class-specific distance threshold. This threshold is calculated from dataset statistics of maximum speed per class, following (zaech2022learnable).
@@ 3.4 Feature Matching and Update
The model is trained with a composite loss. For association, we use the Focal Loss (Lin2017FocalLF) ($\mathcal{L}_{a}$) with $\alpha=-1,\gamma=1$. For position regression, we use a smooth $\ell_{1}$ loss ($\mathcal{L}_{p}$). The total loss is:
@@ 4.1 Experiment Setup
We conduct experiments on two large-scale autonomous driving benchmarks. The Waymo Open Dataset (sun2020scalability) provides 1150 sequences (798 training, 202 validation, 150 testing) of 20-second driving data with 3D labels for three classes: Vehicle, Pedestrian, and Cyclist. The nuScenes dataset (Caesar2019nuScenesAM), our primary benchmark, comprises 1000 sequences of similar length. It is distinguished by a richer sensor suite (32-beam LiDAR, RADAR, six cameras) and provides 3D tracking annotations at 2Hz for seven object classes, making it ideal for leveraging multi-modal data.
@@
We follow the official nuScenes tracking benchmark protocol for evaluation. The primary metrics are Average Multi-Object Tracking Accuracy (AMOTA) and Average Multi-Object Tracking Precision (AMOTP) (Weng20193DMT). We also report secondary metrics from CLEAR MOT (Bernardin2008EvaluatingMO), including Multi-Object Tracking Accuracy (MOTA), ID Switches (IDS).
@@
We construct the training dataset using outputs from CenterPoint (yin2021center). Following 3DMOTFormer (Ding2023MOTFormer), we trained our tracker for 16 epochs on randomly sampled mini-sequences of length TT=10. For the memory buffer, we set the maximum temporal length $T_{max}$=7 and select the top $k=3$ spatially dependent objects to capture the historical information for each tracklet. In the cue-consistent attention module, we set top $k=3$ to perform the cue-consistent cross attention. For optimization, we use the AdamW (loshchilov2017decoupled) optimizer with an initial learning rate of $2e^{-4}$ and a cosine decay schedule with power set to 0.8. All experiments are trained on eight NVIDIA 4090 GPUs with a batch size of one per GPU.
@@ 4.2 Main Results
To validate our approach, we compare DSC-Track against state-of-the-art methods on the nuScenes and Waymo benchmarks. On nuScenes, we compare against two categories of baselines: methods leveraging additional 3D appearance cues like ShaSTA (sadjadpour2023shasta) and NEBP (liang2023neural), and those without, including strong competitors like 3DMOTFormer (Ding2023MOTFormer), PolarMOT (kim2022polarmot), ImmortalTracker (wang2021immortal), SimpleTrack (pang2022simpletrack), and CBMOT (benbarka2021score). On Waymo, we compare against recent strong methods such as SpOT (stearns2022spot) and TrajectoryFormer (chen2023trajectoryformer). For fair comparison, all methods are built upon detections from the widely-used CenterPoint (yin2021center) detector.
@@
This strong performance is mirrored on the nuScenes validation set (Table 2), where DSC-Track again sets a new state-of-the-art with 73.2% AMOTA, surpassing 3DMOTFormer by 2.0%. To evaluate generalizability, we also test on the Waymo Open Dataset (Table 3), where our method excels at tracking stability, securing the best or tied-for-best IDS scores across all major categories: Vehicle, Pedestrian, and Cyclist.